# Rethinking Impersonation and Dodging Attacks on Face Recognition Systems

### v3

@@ 4.4 Analytical Study
The Trade-off Between the Impersonation Attacks and Dodging Attacks: Owing to the inherent conflict during the optimization process of impersonation and dodging losses in the black-box setting, there exists a trade-off between impersonation and dodging performance. We craft adversarial face examples using Lagrangian attack and our proposed Adv-Pruning on LFW based on DI. The average black-box results are demonstrated in the left plot of Fig. 7.
@@
Figure 7: (a) The trade-off between the ASR (%) of impersonation attack and dodging attack of adversarial examples. (b) The dodging ASR (%) in different sparsity ratios.
@@
5 Conclusion
@@
In this paper, we reveal the universality of multi-identity samples among adversarial face examples crafted by previous impersonation attacks and the success of an impersonation attack may not necessarily imply the success of dodging attacks on FR systems in the black-box setting. In order to improve dodging performance without compromising impersonation performance, we proposed a novel attack, namely Adv-Pruning. Adv-Pruning comprises Priming, Pruning, and Restoration Stages. Leveraging our proposed Adversarial Priority Quantification, we identify less prioritized adversarial perturbations with minimal impact on absolute model output variances. Through our proposed Biased Gradient Adaptation, biased gradient perturbations are applied to the sparsified regions, adapting adversarial face examples to a space favoring dodging attacks. Extensive experiments demonstrate the effectiveness of our proposed method.
Appendix A The Pseudo-code of Adversarial Pruning Attack Method
The pseudo-code of our proposed method based on the Lagrangian attack is illustrated in Algorithm 1.
Algorithm 1 Adversarial Pruning Attack Based on Lagrangian
0:  Negative face image pair $\{x^{s},x^{t}\}$, the number of the iterations of Priming stage $n$, the maximum number of iterations $m$, maximum allowable perturbation magnitude $\epsilon$, the surrogate FR model $\mathcal{F}$, the step size $\beta$ and $\gamma$.
0:  An adversarial face example $x^{adv}_{m}$
1:  $x^{adv}_{0}=x^{s}$
2:  for $t=1,...,n$ do
3:     $\mathcal{L}^{p}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$triangleright Priming Stage
4:     $x^{adv}_{t}=x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)$
5:     $x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{t}\right)$
6:  end for
7:  $\mathcal{I}=|x^{adv}_{n}-x^{s}|$triangleright Pruning Stage
8:  $\mathcal{Q}={\rm Sort}\left(\Psi\left(\mathcal{I}\right)\right)$
9:  $\mathcal{W}=\mathcal{Q}\left[:\kappa s\right]$
10:  Get $\mathcal{M}$ by Eq. 8.
11:  $x^{adv}_{n}=x^{s}+\left(x^{adv}_{n}-x^{s}\right)\odot\mathcal{M}$
12:  $\mathcal{A}=1-\mathcal{M}$
13:  for $t=n+1,...,m$ do
14:     $\mathcal{L}^{r}=\tilde{\lambda}\mathcal{L}^{i}+\mathcal{L}^{d}$triangleright Restoration Stage
15:     $x^{adv}_{t}=x^{adv}_{t-1}-\gamma{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}\right)$
16:     $p=\mathcal{A}\odot\left(x^{adv}_{t}-x^{adv}_{n}\right)$
17:     $x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{n}+p\right)$
18:  end for
Appendix B The Optimization Objective of the Impersonation and Dodging Attacks
The objective of impersonation attacks in our research is to craft an adversarial example $x^{adv}$ that causes the model $\mathcal{F}^{vct}$ to incorrectly classify it as the target sample $x^{t}$, while simultaneously maintaining a high degree of visual similarity between $x^{adv}$ and the original sample $x^{s}$. To be more specific, the objective can be expressed as follows:
$\begin{gathered}x^{adv}=\mathop{\arg\min}\limits_{x^{adv}}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{t}\right)\right)\right)\\ \text{s.t.}\|x^{adv}-x^{s}\|_{p}\leq\epsilon\end{gathered}$ (13)
where $\mathcal{D}$ refers to a predefined distance metric, while $\epsilon$ specifies the maximum magnitude of permissible perturbation.
In contrast, the dodging attacks introduced in this study aim to prevent the model $\mathcal{F}^{vct}$ from correctly recognizing $x^{adv}$ as the source sample $x^{s}$. This is achieved while maintaining a visually similar appearance between the adversarial example $x^{adv}$ and the original input $x^{s}$. In a manner similar to impersonation attacks, the objective of dodging attacks can be formulated as follows:
$\begin{gathered}x^{adv}=\mathop{\arg\max}\limits_{x^{adv}}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{s}\right)\right)\right)\\ \text{s.t.}\|x^{adv}-x^{s}\|_{p}\leq\epsilon\end{gathered}$ (14)
Appendix C The proof of the supremum of the absolute variance
In this section, we will provide the proof of the supremum of the absolute variance of the output of $\mathcal{F}$ before and after pruning. The absolute variance can be expressed as follows:
$\rho=\|\mathcal{F}\left(x+\delta\right)-\mathcal{F}\left(x\right)\|$ (15)
[…]
Let us revisit an early hypothesis posited by Goodfellow et al. [14], suggesting that the linear nature of modern DNNs, resembling linear models trained on the same dataset, is the underlying cause of adversarial examples and their surprising transferability [15]. Based on the hypothesis, the FR model $\mathcal{F}$ can be represented in the following linear form:
$\mathcal{F}\left(x\right)=\omega x+b$ (16)
where $\omega$ is the weight vector, and $b$ is the bias term.
Hence, the absolute variance of the output of the function $\mathcal{F}$ can be articulated as:
$\begin{split}\rho&=\|\mathcal{F}\left(x+\delta\right)-\mathcal{F}\left(x\right)\|\\ &=\|\omega\left(x+\delta\right)+b-\left(\omega x+b\right)\|\\ &=\|\omega\delta\|\leq\|\omega\|\|\delta\|=\|\omega\|\mathcal{I}\\ \end{split}$ (17)
The supremum of the absolute variance is given by $\lVert\omega\rVert\mathcal{I}$. Within the context of gradient-based adversarial attacks, the parameters of the neural networks are typically held constant. Consequently, the vector $\omega$ can be treated as a constant. From this, we can deduce that the supremum of the absolute variance in the output of $\mathcal{F}$ is directly proportional to the priority measure $\mathcal{I}$. It follows that adversarial perturbations of greater magnitude are more prioritized, as they exert a more pronounced influence on the output of the face recognition model $\mathcal{F}$.
Appendix D The Calculation Method of the Attack Success Rate
Due to the different objectives of the impersonation and dodging attacks, the calculation methods for the ASR also vary between the two attack types. When it comes to impersonation attacks on FR, the ASR can be computed as:
${\rm ASR}^{i}=\frac{\sum_{i=1}^{N_{p}}\mathds{1}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{t}\right)\right)<t^{i}\right)}{N_{p}}$ (18)
where $N_{p}$ refers to the total number of face pairs and $t^{i}$ represents the impersonation attack threshold. For dodging attacks on FR systems, the ASR can be computed via the following formula:
${\rm ASR}^{d}=\frac{\sum_{i=1}^{N_{p}}\mathds{1}\left(\mathcal{D}\left(\mathcal{F}^{vct}\left(x^{adv}\right),\mathcal{F}^{vct}\left(x^{s}\right)\right)>t^{d}\right)}{N_{p}}$ (19)
where the value of $t^{d}$ represents the threshold for dodging attacks.
Appendix E The Calculation Method of the Multi-identity Sample Ratio
Multi-identity Sample Ratio (MSR) is the proportion of multi-identity samples that can be recognized as both the attacker and victim identities in the adversarial face examples capable of executing successful impersonation attacks. The detailed calculation method of MSR be expressed as:
${\rm MSR}=\frac{|\mathcal{S}_{1}\cap\mathcal{S}_{2}|}{|\mathcal{S}_{2}|}$ (20)
where $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are the set of the adversarial examples that are identified as the attacker and victim identities, respectively. Therefore, the $\mathcal{S}_{2}$ is the set of adversarial face examples that can perform a successful impersonation attack.
Table 4: Comparisons of ASR results (%) of dodging attack and impersonation attacks on the LFW dataset using the negative cosine similarity loss. The models in the second row are the victim models.
ASRd ASRi
IR152 FaceNet MF
Baseline 2.2 16.1 99.4 25.4
Baseline + Lagrangian 4.0 26.5 100.0 26.0
Baseline + ours 8.1 39.7 100.0 26.2
Appendix F Results on Negative Cosine Similarity Loss
Apart from the normalized $L_{2}$ distance loss (i.e. Eq. 1 and Eq. 2), the negative cosine similarity loss is also a commonly used loss function in adversarial attacks on FR systems. In this section, we will assess the effectiveness of our proposed method regarding the negative cosine similarity loss.
For impersonation attacks, the negative cosine similarity loss can be expressed as:
$\mathcal{L}^{i}_{c}=1-\cos\left(\mathcal{F}\left(x^{adv}\right),\mathcal{F}\left(x^{t}\right)\right)$ (21)
where $\cos\left(e,e^{\prime}\right)$ is the function to calculate the cosine similarity between $e$ and $e^{\prime}$.
Similarly, the negative cosine similarity loss for dodging attacks can be formulated as follows:
$\mathcal{L}^{d}_{c}=1+\cos\left(\mathcal{F}\left(x^{adv}\right),\mathcal{F}\left(x^{s}\right)\right)$ (22)
To evaluate the effectiveness of our proposed method on the negative cosine similarity loss, we use DI as the Baseline to craft adversarial face examples using the negative cosine similarity loss $\mathcal{L}^{i}_{c}$ and $\mathcal{L}^{d}_{c}$ as $\mathcal{L}^{i}$ and $\mathcal{L}^{d}$, respectively. The performance is demonstrated in Table 4.
Table 4 demonstrates that the performance of both impersonation and dodging improves after integrating our proposed attack method into the Baseline attack. This further highlights the effectiveness of our method regarding the negative cosine similarity loss.
Table 5: The results (%) on the LFW dataset using MF as the surrogate model. The models on the first row are the victim models.
ASRd ASRi
IR152 FaceNet MF
SA 99.4 45.9 2.1 44.5
SA + Lagrangian 100.0 49.2 3.4 44.8
SA + ours 100.0 54.9 4.7 46.4
Appendix G Results on Sibling-Attack
Sibling-Attack [29] (abbreviated as SA) is a cutting-edge adversarial attack on FR that demonstrates considerable effectiveness in enhancing the transferability of adversarial attacks on FR. SA employs a surrogate model with a two-branch multi-task neural network, which differs from the surrogate models discussed in Section 4.2. Consequently, we exclusively consider SA for comparison with our method.
We aligned the attack setting with the official SA implementation and integrated our proposed Adv-Pruning with SA. The FFHQ dataset, provided by the official SA page [^1] https://github.com/Tencent/TFace/tree/Adv-Attack-Defense/security/tasks/Adv-Attack-Defense/Sibling-Attack, is chosen as the dataset for generating adversarial face examples. The results are depicted in Table 5.
Table 5 demonstrates that the incorporation of our proposed method results in enhancements in both dodging and impersonation performance. This further demonstrates the effectiveness of our proposed attack method in enhancing the dodging attack performance without compromising the impersonation attack performance.
Table 6: The results (%) on the LFW dataset using MF as the surrogate model. The models on the first row are the victim models.
Attack Metric IR152 FaceNet MF
DI MSR 95.6 75.7 0.8
ASRi 18.4 32.9 100.0
ASRd 2.2 18.2 99.2
VMI MSR 97.0 88.6 0.3
ASRi 13.6 20.2 100.0
ASRd 1.0 8.4 99.7
SSA MSR 98.6 92.8 1.5
ASRi 13.8 19.6 100.0
ASRd 0.7 5.9 98.5
DFANet MSR 98.6 96.6 0.4
ASRi 7.0 11.9 100.0
ASRd 0.2 1.5 99.6
SIA MSR 97.4 83.9 1.6
ASRi 15.7 26.7 100.0
ASRd 1.0 10.6 98.4
BSR MSR 98.1 96.8 15.1
ASRi 5.4 9.5 100.0
ASRd 0.4 3.7 84.9
BPFA MSR 96.2 92.1 2.3
ASRi 15.8 17.8 100
ASRd 0.9 4.6 97.7
Appendix H Results on MSR of Previous Impersonation Attacks
Following the same experimental setting of Section 4.4, we conduct experiments on various previous impersonation attacks. The results are shown in Table 6.
As indicated by Table 6, the majority of the crafted adversarial face examples are multi-identity samples in the black-box setting. This observation suggests that the majority of adversarial face examples, which are generated through the previous impersonation attacks, fail to achieve a successful dodging attack in the black-box setting.

### v4

@@ 4.4. Analytical Study
Owing to the inherent conflict during the optimization process of impersonation and dodging losses in the black-box setting, there exists a trade-off between impersonation and dodging performance. We craft adversarial face examples using Lagrangian attack and our proposed Adv-Pruning on LFW based on DI. The average black-box results are demonstrated in the left plot of Figure 7.
@@
Figure 7. (a) The trade-off between the ASR (%) of impersonation attack and dodging attack of adversarial examples. (b) The dodging ASR (%) in different sparsity ratios.
@@
The analytical study of multi-identity samples among the benign face images and universality of multi-identity samples among adversarial face examples are in the supplementary.
5. Conclusion
@@
In this paper, we delve into the issue of multi-identity samples among adversarial face examples. Our research reveals the universality of multi-identity samples among adversarial face examples crafted by previous impersonation attacks and the success of an impersonation attack may not necessarily imply the success of dodging attacks on FR systems in the black-box setting. In order to improve dodging performance without compromising impersonation performance, we proposed a novel attack, namely Adv-Pruning. Adv-Pruning comprises Priming, Pruning, and Restoration Stages. Leveraging our proposed Adversarial Priority Quantification, we identify less prioritized adversarial perturbations with minimal impact on absolute model output variances. Through our proposed Biased Gradient Adaptation, biased gradient perturbations are applied to the sparsified regions, adapting adversarial face examples to a space favoring dodging attacks. Extensive experiments demonstrate the effectiveness of our proposed method.
Acknowledgements.
This work was supported in part by the Natural Science Foundation of China under Grant 61972169,62372203 and 62302186, in part by the National Key Research and Development Program of China (2022YFB2601802), in part by the Major Scientific and Technological Project of Hubei Province (2022BAA046, 2022BAA042), in part by the Knowledge Innovation Program of Wuhan-Basic Research, in part by China Postdoctoral Science Foundation 2022M711251.